OpenStudy (anonymous):

5/4x + 1/8x= 3/8+x the solution is x= ??

OpenStudy (anonymous):

5 / 4x + 1 / 8x = 3 /8 + x 10 / 8 x + 1 / 8x = 3 / 8 + x 11 / 8x = 3 / 8 + x 11 / 8x - x = 3 / 8 3 / 8x = 3 / 8 3 / 3 = 8 x / 8 x = 1

OpenStudy (anonymous):

thank you :)